theyadoreshayyy:

A square prism and a cylinder have the same height. The area of the cross-section of the square prism is 157 square units, and the area of the cross-section of the cylinder is 50π square units. Based on this information, which argument can be made? The volume of the square prism is one third the volume of the cylinder. The volume of the square prism is half the volume of the cylinder. The volume of the square prism is equal to the volume of the cylinder. The volume of the square prism is twice the volume of the cylinder.

theyadoreshayyy:

@snowflake0531

snowflake0531:

@AZ @florisalreadytaken dun dun dun~

Florisalreadytaken:

we know that the formula to find volume for a square prism is: \[ V=a^2\times h \Rightarrow V=b(base) \times h \] for the cylinder it would be: \[ V= \pi r^2 \times h \Rightarrow V=b(base) \times h \] "The area of the `cross-section of the square prism` is `157` square units" that basically means that the base is 157 "area of the `cross-section of the cylinder` is `50π` square units." that also means that the area of the base of that cylinder is \( 50 \pi \) which is equal to 157.1, nearly the same as the prism. that said, what do you think is the right answer?

Florisalreadytaken:

(they have share the same height right, thats why i didnt mention it)

theyadoreshayyy:

is it c ?

Florisalreadytaken:

yes.

theyadoreshayyy:

thank you